# Filtering with Confidence: When Data Augmentation Meets Conformal Prediction

### v1

@@ 3.3 Learning to recognize quality outputs on $\mathcal{D}_{\text{train}}$
In other words, the gold-standard is the population average of the observed surrogate, and conversely, $\tilde{\mathcal{A}}$ can be viewed as a specific realization of a random variable, centered at $\mathcal{A}$. For instance, in text data, embedding-based similarity metrics such as cosine similarity computed from BERT embeddings are widely used to capture semantic coherence Devlin et al. (2018); Zhang et al. (2020). In image data, similarity measures based on CLIP scores Radford et al. (2021) are effective for capturing both semantic alignment and stylistic similarity. These metrics typically compare each generation directly against its original sample, which can be viewed as a realization from the underlying distribution $h^{\star}$.
We propose reducing the variability of the surrogate $\tilde{\mathcal{A}}$ by applying a regression-based strategy that leverages similar samples to approximate the underlying expectation. By smoothing over similar samples, this learned approximation is expected to more closely reflect the ground-truth measure $\mathcal{A}$.
@@ 4 Experiments
To highlight the efficacy of our method, we propose three case studies: (a) a data enrichment setting, (b) an imbalanced classification setting, and (c) a very low-data regime with generations of heterogeneous quality. Our examples span different data types, from text, to images, to tabular data.
@@ 4.1 Prediction with LLM-augmented training data
We study our data augmentation pipeline for clinical text classification, focusing on mapping symptom descriptions to medical diagnoses (Gretel AI, 2024). The dataset $(X_{i},Y_{i})$} consists of 853 training samples and 212 test samples, where $X_{i}$ denotes a symptom description and $Y_{i}$ is one of 22 possible diagnoses. Each training example $(X_{i},Y_{i})$ is augmented using a generative language model (GPT-4.1 nano (OpenAI, 2023)), which extends the original symptom description $X_{i}$ with five additional sentences $(G_{ik})_{k=1}^{5}$. From these extensions, we generate new samples that inherit the original label, yielding a total of 4,265 synthetic observations $\{(G_{ik})_{k=1}^{5}:i=1,2,\cdots,853\}$.
To ensure output quality, we employ a two-stage evaluation strategy. First, a random subset of 500 generations, derived from 100 symptom descriptions, is evaluated with a high-accuracy model $\mathcal{A}$ (Gemini-2.5-pro (Gemini Team, Google DeepMind, 2023)), forming the calibration set $\mathcal{D}_{\text{calib}}$. Next, all augmented samples are scored using a faster, lower-cost surrogate model $\hat{\mathcal{A}}$ (Gemini-2.5-flash (Gemini Team, Google DeepMind, 2023)). Both models assign a score in $[0,1]$, with $0.5$ as the retention threshold (see the detailed prompt in the supplement). This design reflects a practical labeling scenario in which reliable annotations are costly, whereas approximate labels can be obtained inexpensively. Let $(A_{ik})_{i\in\mathcal{D}_{\text{calib}}}$ denote the Gemini-pro scores and $(\hat{A}_{ik})_{i=1}^{853}$ denote the Gemini-flash scores.
We then apply our calibration step. For each $(X_{i},(G_{ik})_{k=1}^{K},Y_{i})\in\mathcal{D}_{\text{calib}}$, a non-conformity score is defined as the minimum threshold that guarantees all selected sentences achieve a pro-score above $0.5$, so that: $S(\hat{\mathbf{A}}_{i},\mathbf{A}_{i})=\inf\left\{\tau:\big|\{G_{ik}:1\leq k\leq 5,\,\hat{A}_{ik}\geq\tau,\,A_{ik}<0.5\}\big|\leq 1\right\}.$
@@
To evaluate performance, we fine-tune a diagnostic classifier (distilbert-base-uncased (Devlin et al., 2018)) using LoRA (Hu et al., 2022). Each training iteration consists of 100 fixed high-confidence documents (selected by the pro-scores) and 400 additional documents sampled under one of the following filtering schemes: (1) No augmentation; (2) No filtering; (3) Filtering by $\hat{A}_{ik}$ only (threshold = 0.5); (4) Hybrid filtering (using $A_{ik}$ for $\mathcal{D}_{\text{calib}}$ and $\hat{A}_{ik}$ with threshold 0.5 for the remainder); (5) CondCP-based filtering on $\mathcal{D}_{\text{aug}}$ (using $A_{ik}$ for $\mathcal{D}_{\text{calib}}$ ).
Performance, averaged across 20 trials, is reported in Figure 2, with evaluation consistently conducted on the held-out test set. We also report the results of experiments performed in an identical manner on topic prediction (predicting the topic of statistical abstracts downloaded from arXiv with 5 possible categories) and sentiment analysis (predicting one of 6 emotions on a dataset of Twitter messages — see details in the Appendix D). Overall, across these three datasets, our CondCP filter improves the precision, recall, and F1-score by up to 3% over the unaugmented baseline, and substantially improves upon the unfiltered baseline. We note that in the diagnosis task, the unfiltered augmentation outperforms the unaugmented baseline, but this advantage does not hold for the abstract and Twitter datasets, suggesting that including all generations can be detrimental when low-quality samples are present. In contrast, the CondCP filter achieves the best performance across all metrics and tasks.
Figure 2: Evaluation of different data augmentation methods on diagnosis prediction, abstract topic prediction, and Twitter message sentiment prediction. Results are averaged over 20 replicates. Error bars indicate the interquartile range, with centers representing the median and boundaries corresponding to the first and third quartiles.

### v2

@@ 3.3 Learning to recognize quality outputs on $\mathcal{D}_{\text{train}}$
In other words, the gold-standard is the population average of the observed surrogate, and conversely, $\tilde{\mathcal{A}}$ can be viewed as a specific realization of a random variable, centered at $\mathcal{A}$. For instance, in text data, embedding-based similarity metrics such as cosine similarity computed from BERT embeddings are widely used to capture semantic coherence (Devlin et al., 2019; Zhang et al., 2019). In image data, similarity measures based on CLIP scores (Radford et al., 2021) are effective for capturing both semantic alignment and stylistic similarity. These metrics typically compare each generation directly against its original, which can be viewed as a realization from the underlying distribution $h^{\star}$.
We propose reducing the variability of the surrogate $\tilde{\mathcal{A}}$ by applying a regression-based strategy that leverages similar samples to approximate the underlying expectation. By smoothing over similar samples, this learned approximation is expected to reflect more closely the true measure $\mathcal{A}$.
@@ 4 Experiments
To highlight the efficacy of our method, we propose three case studies: (a) a data enrichment setting, (b) an imbalanced classification setting, and (c) a low-data regime with generations of heterogeneous quality. Our examples span different data types, from text, to images, to tabular data.
@@ 4.1 Prediction with LLM-augmented training data
We study our data augmentation pipeline for clinical text classification, focusing on mapping symptom descriptions to medical diagnoses (Gretel AI, 2024). The dataset $(X_{i},Y_{i})$} consists of 853 training samples and 212 test samples, where $X_{i}$ denotes a symptom description and $Y_{i}$ is one of 22 possible diagnoses. Each training example $(X_{i},Y_{i})$ is augmented using a generative language model (GPT-4.1 nano (Achiam et al., 2023)), which extends the original symptom description $X_{i}$ by generating five additional sentences $(G_{ik})_{k=1}^{5}$ using a temperature parameter $\tau=1.5$, to encourage the creation of varied content. (In Appendix D.8, we report results for a temperature of $\tau=0.5$ to compare the results, and show in that case that the augmentation has limited efficacy). From these extensions, we generate new samples that inherit the original label, yielding a total of 4,265 synthetic observations $\{(G_{ik})_{k=1}^{5}:i=1,2,\cdots,853\}$.
To ensure output quality, we employ a two-stage evaluation strategy. First, a random subset of 500 generations, derived from 100 symptom descriptions, is evaluated with a high-accuracy model $\mathcal{A}$ (Gemini-2.5-pro (Comanici et al., 2025)), forming the calibration set $\mathcal{D}_{\text{calib}}$. Next, all augmented samples are scored using a faster, lower-cost surrogate model $\hat{\mathcal{A}}$ (Gemini-2.5-flash (Comanici et al., 2025)). Both models assign a score in $[0,1]$, with $0.5$ as the retention threshold (see the detailed prompt in the supplement). This design reflects a practical labeling scenario in which reliable annotations are costly, whereas approximate labels can be obtained inexpensively. Let $(A_{ik})_{i\in\mathcal{D}_{\text{calib}}}$ denote the Gemini-pro scores and $(\hat{A}_{ik})_{i=1}^{853}$ denote the Gemini-flash scores.
We then apply our calibration step. For each $(X_{i},(G_{ik})_{k=1}^{K},Y_{i})\in\mathcal{D}_{\text{calib}}$, a non-conformity score is defined as the minimum threshold that guarantees all selected sentences achieve a pro-score above $0.5$, so that:
$\quad S(\hat{\mathbf{A}}_{i},\mathbf{A}_{i})$
$=\inf\left\{\tau:\big|\{G_{ik}:1\leq k\leq 5,\,\hat{A}_{ik}\geq\tau,\,A_{ik}<0.5\}\big|\leq 1\right\}.$
@@
Prediction Performance.
To evaluate performance, we fine-tune a diagnostic classifier (distilbert-base-uncased (Devlin et al., 2019)) using LoRA (Hu et al., 2022). Each training iteration consists of 100 fixed high-confidence documents (selected by the pro-scores) and 400 additional documents sampled under one of the following filtering schemes: (1) No augmentation; (2) No filtering; (3) Filtering by $\hat{A}_{ik}$ only (threshold = 0.5); (4) Hybrid filtering (using $A_{ik}$ for $\mathcal{D}_{\text{calib}}$ and $\hat{A}_{ik}$ with threshold 0.5 for the remainder); (5) CondCP-based filtering on $\mathcal{D}_{\text{aug}}$ (using $A_{ik}$ for $\mathcal{D}_{\text{calib}}$).
Performance, averaged across 20 trials, is reported in Figure 2(a), with evaluation consistently conducted on the held-out test set. We also report the results of experiments performed in an identical manner on topic prediction (predicting the topic of statistical abstracts downloaded from arXiv with 5 possible categories) and sentiment analysis (predicting one of 6 emotions on a dataset of Twitter messages — see details in the Appendix D). Overall, across these three datasets, our CondCP filter improves the precision, recall, and F1-score by up to 3 percentage points over the unaugmented baseline, and substantially improves upon the unfiltered baseline. We note that in the diagnosis task, the unfiltered augmentation outperforms the unaugmented baseline, but this advantage does not hold for the abstract and Twitter datasets, suggesting that including all generations can be detrimental when low-quality samples are present. In contrast, the CondCP filter achieves the best performance across all metrics and tasks.
Out-Of-Domain Performance.
To evaluate robustness and generalization under domain shift, we report the performance of different data augmentation frameworks on out-of-domain tasks in Figure 2(b). Specifically, we fine-tune the BERT classifier on statistical abstracts and evaluate it on abstracts covering the same topics but drawn from different domains, such as statistical computation versus physical computation and computer science numerical analysis (see Appendix D.4). The CondCP filter achieves the best performance, suggesting our data augmentation strategy improves the model’s ability to capture task-relevant semantics rather than domain-specific artifacts.
Figure 2: (a) Performance of different data augmentation methods on three tasks: diagnosis prediction, abstract topic prediction, and Twitter message sentiment analysis. Results are averaged over 20 replicates. Error bars denote the interquartile range (IQR), with centers representing the median and boundaries corresponding to the first and third quartiles. (b) Out-of-domain (OOD) evaluation on abstract topic prediction: the classifier is trained on statistical abstracts and tested on abstracts from different domains.